OpenStudy (anonymous):

The box plots show male and female grades in a U.S. History class. two bar graphs shown. The top one is labeled Male. Minimum at 70, Q1 at 82, median at 88, Q3 at 95. The bottom bar graph is labeled Female. Minimum at 75, Q1 at 78, Q3 at 90, maximum at 92. Which of the following best describes the information about the interquartile ranges? The female interquartile range cannot be known because only three quartiles are showing. The male interquartile range cannot be known because the fourth quartile is missing. The male interquartile range is larger due to the size of the first

OpenStudy (anonymous):

I said its A is this correct ?

OpenStudy (kirbykirby):

The interquartile range is quantity Q3 - Q1 We know these values for both males and females. For males: Q3 - Q1 = 95 - 82 = 13 For females: Q3 - Q1 = 90 - 78 = 12